A rod with rectangular cross section oscillates about a horizontal axis passig through one of its ends and it behave like a second pendulum. Determine it`s length

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
T = 2*pie sqrt(I/mgd) - It is Rigid body 
By parallel axis theorem 
I = ML^2/12 + M(L/2)^2 = Ml^2/3 
d = l/2 
T = 2*pie sqrt(2ML^2/3MgL) 
= 3/2 *1 = 1.5m